OpenStudy (anonymous):

Find the difference. 14x - (x2 + 8x) =

OpenStudy (anonymous):

my best answer is 4x

OpenStudy (anonymous):

in the parenthesis you add both numbers which is 10x, then you subtract 14x from 10x which will give you 4x :)

OpenStudy (anonymous):

\[14x-(x^{2}+8x)=\]\[14x-x^{2}-8x=\]\[-x^{2}-6x.\]

OpenStudy (anonymous):

well can you tell me how you got that? hmm ok i see were you get the anwser but tbh im still lost @-@

OpenStudy (anonymous):

\[-x^2-6x\]

OpenStudy (anonymous):

look at my first post :)

OpenStudy (anonymous):

can you use this as an exampel to show me ??? (6x3 + 18x2) - 3x2 =

OpenStudy (anonymous):

oh mines wrong, i did't see the x being squared lol take their answers ^.^

OpenStudy (anonymous):

\[(6x^3 + 18x^2) - 3x^2 =\]\[6x^3 + \color{red} { 18x^2 } \color{red} { - 3x^2 } =\]\[6x^3 + \color{red} { 15x^2 } =\]\[6x^3 + \color{black} { 15x^2 } \]

OpenStudy (anonymous):

ok i have to try that ! im still lost but i see what you guys are goiing kinda thanks!!!

OpenStudy (anonymous):

(16x3 + 2x2) - 4x3 = ima try this one can you tell me if im right??

OpenStudy (anonymous):

sure, go ahead :)

OpenStudy (anonymous):

oh and the two 3 and two 2s are expomets

OpenStudy (anonymous):

yes... so do you want to do it, or should I ?

OpenStudy (anonymous):

i got 13x^3+2x^2??? is that right???

OpenStudy (anonymous):

close, but it would be 12x^3, b/c 16-4=12, not 13.

OpenStudy (anonymous):

ohh haha simpel mistake

OpenStudy (anonymous):

yes, but I am sure that you are capable of doing this math .

OpenStudy (anonymous):

thank you!! you helped alot!! i have one more question tho... what do you do it there is a neagtive on the outside of a parinthasy like this equation? 7x - (25x2 + 12x) =

OpenStudy (anonymous):

\[7x - (25x^2 + 12x) =\]\[7x - 25x^2 - 12x \] see what I am doing in this step ? Now.... \[\color{black} { 7x } - 25x^2 \color{black} { - 12x } \] \[\color{red} { 7x } - 25x^2 \color{red} { - 12x } \] \[\color{red} { 7x-12x } - 25x^2\] \[\color{red} { -5x } - 25x^2\] \[ - 25x^2\color{black} { -5x }\]

OpenStudy (anonymous):

so basicaly the () dont need to be there?

OpenStudy (anonymous):

well, when you have -(a+b) we simplify it into -a-b

OpenStudy (anonymous):

ohh ok i kinda understand

OpenStudy (anonymous):

Good ☻